is cos x = 0.25, find sin x
i have looked up tutorials but i am a bit confused.

Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

cos x^2+sin x^2=1, a useful identity.
so 0.25^2 +sin x^2=1
sin x^2=0.9375, since 0.25^2=0.0625
sin x=0.9682
-
can check with tables
cos x=0.25 when x=75.52 degrees, using the inv cos function. Then with the degrees/radians, look at the sine function.
sin 75.52 deg.=0.9682

Answer by MathTherapy(10552)

is cos x = 0.25, find sin x
i have looked up tutorials but i am a bit confused.